# Can Classic GNNs Be Strong Baselines for Graph-level Tasks? Simple Architectures Meet Excellence

### v2

@@ Abstract
[…]
@@ 3.5 Feed-Forward Network
GCN+ 0.076 $\pm$ 0.009 79.3$\downarrow$ 98.382 $\pm$ 0.095 8.5$\uparrow$ 69.824 $\pm$ 0.413 25.4$\uparrow$ 87.021 $\pm$ 0.095 21.1$\uparrow$ 77.109 $\pm$ 0.872 12.6$\uparrow$
GIN 0.526 $\pm$ 0.051 96.485 $\pm$ 0.252 $5$$\pm$ 1.527 85.387 $\pm$ 0.136 64.716 $\pm$ 1.553
GIN+ 0.065 $\pm$ 0.004 87.6$\downarrow$ 98.285 $\pm$ 0.103 1.9$\uparrow$ 69.592 $\pm$ 0.287 25.9$\uparrow$ 86.842 $\pm$ 0.048 1.7$\uparrow$ 74.794 $\pm$ 0.213 15.6$\uparrow$
@@
GatedGCN+ 0.077 $\pm$ 0.005 72.7$\downarrow$ 98.712 $\pm$ 0.137 1.4$\uparrow$ 77.218 $\pm$ 0.381 14.7$\uparrow$ 87.029 $\pm$ 0.037 1.7$\uparrow$ 79.128 $\pm$ 0.235 7.1$\uparrow$
@@ 3.6 Positional Encoding
[…]
@@ 4 Assessment: Experimental Setup
GatedGCN+ 0.7006 $\pm$ 0.0033 3.6$\uparrow$ 0.2431 $\pm$ 0.0020 1.9$\downarrow$ 0.4263 $\pm$ 0.0057 9.9$\uparrow$ 0.3802 $\pm$ 0.0015 30.1$\uparrow$ 0.9460 $\pm$ 0.0057 2.6$\uparrow$
@@ 5.2 Ablation Study
Ablating FFN leads to substantial performance declines for GIN+ and GCN+ across almost all datasets, highlighting its essential role in graph-level tasks. Notably, on MNIST, removing FNN leads to an 88% relative accuracy drop for GIN+. This is likely because the architectures of GIN+ and GCN+ rely heavily on FFN for learning complex node feature representations. In contrast, GatedGCN+ uses gating mechanisms to adaptively adjust the importance of neighboring nodes’ information, reducing the need for additional feature transformations. The only exceptions are observed in the peptides datasets, where FFN is not used in all three models. This may be due to the shallow GNN architecture, where complex feature transformations are less necessary.
@@ 6 Conclusion
This study highlights the often-overlooked potential of classic GNNs in tacking graph-level tasks. By integrating six widely used techniques into a unified GNN+ framework, we enhance three classic GNNs (GCN, GIN, and GatedGCN) for graph-level tasks. Evaluations on 14 benchmark datasets reveal that, these classic GNNs match or outperform SOTA GTs, while also demonstrating greater efficiency. These findings challenge the prevailing belief that GTs are inherently superior, reaffirming the capability of simple GNN structures as powerful models.
@@ Acknowledgments
We sincerely thank the anonymous area chair for their thoughtful evaluation and strong support, which gave us the opportunity to share this work with the community. We are also grateful to all anonymous reviewers for their constructive and insightful feedback. We sincerely thank Yiwen Sun for her invaluable help with refining the writing and to Lu Fan for her help with creating Figure 1.
@@ C.1 Clarification on GNN+ and Classic GNNs
C.2 Related Works
Our work relates closely to recent efforts in benchmarking GNNs. As node-level benchmarks have received significantly more attention, several studies such as (Shchur et al., 2018; Wang et al., 2021; Lv et al., 2021) have examined the reliability of existing evaluation protocols, revealing that many newly proposed architectures fail to outperform simple baselines under fair conditions. Building on this line of inquiry, Luo et al. (2024b) demonstrated that classic GNNs can match or even surpass SOTA models across 18 widely used node classification datasets.
C.3 Limitations and Future Work
While our study presents strong empirical evidence that enhanced classic GNNs can serve as strong baselines for graph-level tasks, it remains primarily an empirical benchmarking effort. As such, we do not provide a theoretical analysis of why these classic GNNs can outperform GTs. This omission represents a key limitation of the current work.
A deeper theoretical understanding, for example in terms of expressive power, the mechanisms underlying over-smoothing and over-squashing, and the interplay between architectural components, could offer valuable insights into the design of more effective GNN variants. In future work, we aim to formally characterize the conditions under which GNN+ achieves superior performance and to develop principled analyses of its expressiveness and representational behavior.

### v3

@@ Abstract
Message-passing Graph Neural Networks (GNNs) are often criticized for their limited expressiveness, issues like over-smoothing and over-squashing, and challenges in capturing long-range dependencies. Conversely, Graph Transformers (GTs) are regarded as superior due to their employment of global attention mechanisms, which potentially mitigate these challenges. Literature frequently suggests that GTs outperform GNNs in graph-level tasks, especially for graph classification and regression on small molecular graphs. In this study, we explore the untapped potential of GNNs through an enhanced framework, GNN+, which integrates six widely used techniques: edge feature integration, normalization, dropout, residual connections, feed-forward networks, and positional encoding, to effectively tackle graph-level tasks. We conduct a systematic re-evaluation of three classic GNNs—GCN, GIN, and GatedGCN—enhanced by the GNN+ framework across 14 well-known graph-level datasets. Our results reveal that, contrary to prevailing beliefs, these classic GNNs consistently match or surpass the performance of GTs, securing top-three rankings across all datasets and achieving first place in eight. Furthermore, they demonstrate greater efficiency, running several times faster than GTs on many datasets. This highlights the potential of simple GNN architectures, challenging the notion that complex mechanisms in GTs are essential for superior graph-level performance. Our source code is available at  https://github.com/LUOyk1999/GNNPlus.
@@ 3.5 Feed-Forward Network
[…]
GIN 0.526 $\pm$ 0.051 96.485 $\pm$ 0.252 55.255 $\pm$ 1.527 85.387 $\pm$ 0.136 64.716 $\pm$ 1.553
[…]
@@
[…]
@@ 3.6 Positional Encoding
where $\cdot\|\cdot$] denotes concatenation, $\boldsymbol{x}^{\text{RWSE}}_{v}$ represents the RWSE of node $v$, and $\boldsymbol{W}_{\text{PE}}$ is the trainable weight matrix.
@@ 4 Assessment: Experimental Setup
[…]
@@ 5.2 Ablation Study
Ablating FFN leads to substantial performance declines for GIN+ and GCN+ across almost all datasets, highlighting its essential role in graph-level tasks. Notably, on MNIST, removing FFN leads to an 88% relative accuracy drop for GIN+. This is likely because the architectures of GIN+ and GCN+ rely heavily on FFN for learning complex node feature representations. In contrast, GatedGCN+ uses gating mechanisms to adaptively adjust the importance of neighboring nodes’ information, reducing the need for additional feature transformations. The only exceptions are observed in the peptides datasets, where FFN is not used in all three models. This may be due to the shallow GNN architecture, where complex feature transformations are less necessary.
@@
6 Related Work
Our work relates closely to recent efforts in benchmarking GNNs. As node-level benchmarks have received significantly more attention, several studies such as (Shchur et al., 2018; Wang et al., 2021; Lv et al., 2021; Platonov et al., 2023) have examined the reliability of existing evaluation protocols, revealing that many newly proposed architectures fail to outperform simple baselines under fair conditions. Building on this line of inquiry, Luo et al. (2024b) rigorously showed that, with the same hyperparameter search space, classic GNNs can match or even surpass SOTA models across 18 widely used node classification datasets.
7 Limitations and Future Work
While our study provides strong empirical evidence that enhanced classic GNNs can match or even surpass GTs in graph-level tasks on existing benchmarks, it is important to acknowledge that these findings are derived solely from empirical evaluations and lack a theoretical explanation for the observed advantage. Moreover, as highlighted in recent discussions (Bechler-Speicher et al., 2025), current graph datasets may not adequately capture the complexity and diversity inherent in real-world graph problems. Consequently, although our results demonstrate the competitiveness of classic GNNs under current datasets and evaluation settings, their relative advantage may shift as more challenging and application-driven benchmarks are developed.
A deeper theoretical understanding is needed to elucidate the role of each architectural component within GNN+ and their interactions, particularly regarding the model’s expressiveness, representational behavior, and its capacity to address oversmoothing and oversquashing issues. These directions represent promising avenues for future research.
@@ 8 Conclusion
This study highlights the often-overlooked potential of classic GNNs in addressing graph-level tasks. By integrating six widely used techniques into a unified GNN+ framework, we enhance three classic GNNs (GCN, GIN, and GatedGCN) for graph-level tasks. Comprehensive evaluations on 14 benchmark datasets reveal that these enhanced GNNs can match or even outperform SOTA GTs, while also demonstrating greater computational efficiency. These findings challenge the prevailing belief that GTs are inherently superior, reaffirming the effectiveness of simple GNN structures as powerful models for graph-level tasks.
@@ Acknowledgments
We sincerely thank the anonymous area chair for their thoughtful evaluation and strong support, which gave us the opportunity to share this work with the community. We are also grateful to all anonymous reviewers for their constructive and insightful feedback. We sincerely thank Yiwen Sun for her invaluable help with refining the writing and Lu Fan for her help with creating Figure 1.